OpenStudy (♥ashley♥):

A battery manufacturer randomly tests 500 batteries and finds that 3 are defective. How many defective batteries would be expected in a shipment of 12,000 batteries? A. 120 B. 72 C. 60 D. 36

OpenStudy (♥ashley♥):

I think its B. 72.

undeadknight26 (undeadknight26):

hmm well 12000/500 = ?

OpenStudy (♥ashley♥):

12000/500= 24

undeadknight26 (undeadknight26):

you are correct madam Ashley. because: 12000/500 = 24 24 x 3 = 72

OpenStudy (♥ashley♥):

24 x 3= 72

OpenStudy (anonymous):

I said nothing lol XD

OpenStudy (♥ashley♥):

Thanks everyone!